HUBBLE SPACE TELESCOPE DAILY REPORT #825 PERIOD COVERED: 0800 EST Daily Status Report as of 048/1200Z. 1.0 ACTIVITIES SCHEDULED AND ACCOMPLISHED: 1.1 Completed FOC/96 4071 (UV Studies of Galaxies) The Faint Object Camera (f/96) was used to make three ultraviolet observations to study the structure of the nucleus of galaxy NGC-4278. All three observations were received and displayed. They were reported to look okay. 1.2 Completed WFC 1297 (GTO High Latitude Parallel Program) The Wide Field Camera was used to take four observations in parallel with the prime science observations. All four image were received and displayed in the OSS. They were reported to look okay. 1.3 Completed WFC/PC 4109 (Rapid Internal Monitor) Both the Wide Field Camera and the Planetary Camera were used to make internal flat calibration observations. Both images were received and displayed. They were reported to look okay. 1.4 Completed Two Sets of FOC/96 3591 (Mass Stars in Star Burster Galaxies, FOC Images, Part 1) The Faint Object Camera was used to make two sets of observations of the starbursting galaxy NGC-3991, to study the physics of its massive stars and the processes that drive the formation and evolution of the galaxy. The first set of observations were received and displayed. It was reported to look good, but part of the image fell behind the occulting finger. The second observations was made in fine lock using FGSs #1 and #2. The image was reported to look good. 1.5 Uplinked Updated Ephemeris Tables, Per MOSES 2871 1.6 Completed GHRS/2 4792 (Hydrogen Lyman Alpha Day Glow Emission Line Profiles) The High Resolution Spectrograph was used to make three spectral observations of Jupiter's Lyman Alpha dayglow on its central meridian at the equator and at mid-latitudes. All three observations were received and displayed. They were reported to be okay. 1.7 Completed FOS/RD 3791 (Quasar Absorption Line Survey) The Faint Object spectrograph was used to make observations of quasar PG1254+047 to find out if it is appropriate to make further future observations of this quasar. The spectral observation was reported to be somewhat noisy. 2.0 FLIGHT OPERATIONS SUMMARY: 2.1 Guide Star Acquisitions: NOTE: Only FGS #1 and #3 are being used. Scheduled Acquisitions : 3 Successful : 3 Scheduled Reacquisitions : 5 Successful : 5 Losses of Lock : 0 2.2 FHST Updates: Scheduled : 7 Successful : 6 (FHST Failure 047/2342Z) 2.3 Significant HSTARs: None 2.4 Operations Note: The GSFC Simulation #4 (HST Deploy) is currently under way (048/1300Z - 049/0100Z). 3.0 ACTIVITIES SCHEDULED FOR NEXT REPORTING PERIOD: 3.1 Complete FOS/RD 3483 (Lyman Alpha Emission in Galaxies) 3.2 Complete PC 3603 (Snapshot of Protoplanetary Nebulae) 3.3 Complete PC 3799 (PC Imaging of Gravitation Lenses, WFPC GTO Augmentation) 3.4 Complete FGS 1570 (Extragalactic Astrometry and Astrophysics F10-11, PT 10F2) 4.0 ANOMALY STATUS: 4.1 Significant New Anomalies: None 4.2 Significant Unresolved Anomalies: The FGS #2 anomalous behavior 5.0 SIGNIFICANT FORTHCOMING EVENTS: Nothing at this time. -END- ----- End Included Message -----